the currents I., Iz, and Iz in given Find circuits. lart 182 { 352 2532 2 Ispr 12V 22 IN

Find the currents I1, I2, and I3 in given circuits.

1) 9V, 12V, 25 ohm, 18 ohm, 35 ohm

2) 6V, 12V, 25 ohm, 10 ohm, 15 ohm

3) 6V, 3V, 2 ohm, 10 ohm, 8 ohm, 12 ohm, 6 ohm
Answer #1


Solk SysRa Break ris I In loop AB CDA V - Iz Rz - I R=0 12=35 I2 + 1812 - O и Doop cPE Fc -I, R, - Iz R₂ =O 9=25 I, + 18 Iz - 25 - 1913 Iz + 12 3 - 18 = 12 0.703=2.234 Iz Iz = 0.3146A 7, 25 - x 0.3146 = 0.13365 A 13-12-I, = 0.3146-0.1335 In = 0-1811 A& 10 h + I2 = 2 3 - 1 3 2 II =-0.0234 A I, = 0.3484A 52939 Iz = 0.3484-0.0724 = 0. 27IA 252 Slor 822 12 12 86-(12+8) I, - 612

write in comment box in case of any doubt